# Introducing AI to an Online Petition Platform Changed Outputs but not Outcomes

### v2

@@ Abstract
01_introduction
@@ Results
In our main analysis we use Australia as the control country to estimate causal effects. We confirm that our estimates are robust to multiple geographic regions with a synthetic control analysis that leverages additional control countries (New Zealand, India, and China). The results of the synthetic control analysis are consistent with the difference-in-differences analysis, demonstrating that our causal estimates are consistent across regions and outside of Australia.
@@
We augmented our pre-registered difference-in-differences analysis by studying the impact of AI access on platform homogeneity. The results show an increase in inter-petition homogeneity, as measured through the mean pairwise similarity between petitions [wu2024generativemonoculturelargelanguage, padmakumar2023does]. To measure homogeneity, we first created 3,072-dimension embeddings with the OpenAI text-embedding-3-large model [^3] https://platform.openai.com/docs/models/text-embedding-3-large for all petitions. We then calculated the cosine similarity between all pairs of petitions posted in a given week and country, and took the mean of these pairwise scores to obtain the mean pairwise similarity for that week and country. We used these weekly country-level time series of homogeneity to compare trends in treated and untreated countries through a dynamic difference-in-differences estimation.
@@
In the SI Appendix we provide details of an exploratory predictive model showing that the lexical style of petitions written with access to AI (longer text, higher complexity) was associated that of petitions that had achieved positive outcomes prior to AI tool access. The predictive power of these features declined when users received access to the in-platform AI tool. We consider potential explanations and ramifications in the Discussion section.
@@
To test the consistency of platform participation, we used pre-AI user participation data from treated and untreated countries to fit a Bayesian Structural Time Series (BSTS) model (see details in SI Appendix) [brodersen2015causalimpact]. This BSTS model uses pre-AI long-term trends and seasonal fluctuations to predict a synthetic control for counterfactual participation in treated countries had there been no AI tool launch. This forecast for platform participation closely reflected actual user participation, with actual participation falling within a 95% credible interval of forecast participation. Through comparing the synthetic counterfactual participation against actual platform participation, we do not find evidence of a significant effect of AI access on user participation, in either direction ($\Delta$=-0.45, 95% CI [-2.71, 1.75], p = 0.349).
@@
We further demonstrate that changes to platform composition were not responsible for changes to platform trends by showing that the platform-level trends—shifting petition features and unimproved outcomes—are consistent even when examining trends within repeat users. We focus our analysis here on one lexical feature (word count) and one outcome (comment engagement). The results for the other variables were similar, and are shown in the SI Appendix. We do not conduct a panel difference-in-differences analysis because of the rarity of repeat petition writing; only 7.6% of users in our dataset wrote multiple petitions, which amounted to too few untreated users for a reliable balanced panel.
@@ Discussion
Our work provides a causal analysis of the impact of AI integration into a real-world online platform. Our findings extend prior work on the impact of AI on online platforms at large: while online platforms might be influenced by both off-platform AI and in-platform AI tools, past studies have focused on general-use off-platform AI. As online platforms increasingly integrate AI, it is important to understand the potential impact on platform dynamics. In our case, the staggered roll-out of the Change.org in-platform AI tool allowed us to circumvent the challenges of noisy AI detection, and instead directly measure the impact of in-platform AI access through causal methods. We find that in-platform AI on Change.org shifts petition lexical features and increases site-wide homogeneity. At the same time, the in-platform AI tool did not improve platform outcomes, nor did it seem to affect platform participation. Over time, the in-platform AI integration may create undesirable new harms, such as increased suspicion by content consumers, or reduced effort by content producers.
@@
Why doesn’t AI improve outcomes on Change.org? We have three hypotheses for the lack of change in outcomes: first, that AI might not substantively improve content quality; second, that AI might increase readers’ AI suspicion; and third, that AI-assisted writing might reduce petition-writer ownership and effort. Previous research suggests that when users adopt AI writing tools, their work scores higher in persuasion, creativity, and task performance [schoenegger2025large, Noy2023-sf, dellacquaNavigatingJaggedTechnological2023]. However, AI tools are not guaranteed to augment human performance. In a 2024 meta-review of human-AI collaboration, 15% of reviewed human-AI systems performed worse than humans alone. Dell’Acqua et al. explain the varied success of human-AI collaboration through the ever-expanding “jagged technological frontier” of AI capabilities [dellacquaNavigatingJaggedTechnological2023]. Within this frontier, AI augments human productivity and performance; outside of the frontier, AI performs worse than humans alone. As AI is capable of generating text that contains elements that have been shown to drive petition success, such as persuasion and emotional appeal [HAGEN2016783, chen2019multi, schoenegger2025large], we would expect AI to be capable of generating petitions that achieve strong outcomes. However, the technological capabilities of AI might be insufficient to explain the outcomes of content produced on social platforms, as evaluations are determined by readers’ dynamic, subjective, and social perceptions. Specifically, it is possible that petitions, with their real-world, context-specific missions, fall outside of AI’s capabilities for augmentation because the task requires personal and community-specific knowledge. AI on Change.org might contribute to what others have called “semantic garbage” or “AI slop” [feuerriegel2023research, floridi2020gpt, moller2025impact, tullis2025sifting]: although AI shapes petitions to resemble previous articulate, well-researched petitions, the changes may be superficial, and thus be perceived as less informative or authentic than human-written petitions.
@@
Another way in which readers’ perceptions of petitions might change with in-platform AI access could be AI suspicion. Prior work shows that while people use biased heuristics to detect AI [jakesch2023human], when people suspect AI use, they may perceive the content as less trustworthy and assign lower evaluations to the writers [kadoma2025generative, reif2025socialpenalty, jakesch2019ai, hohenstein2023artificial]. If the AI-written petitions induce higher reader suspicion, outcomes are indeed likely to be affected. This hypothesis is difficult to test with our data since human perceptions of AI assistance may have shifted considerably in the year since our study’s petitions were posted [duan2024understanding, matatov2024examining]. On the other hand, there are reasons to discount the “AI suspicion” hypothesis in this case. Petition readers were not primed to suspect AI use on Change.org; during the ten-week period for which we estimate the causal impact of AI, Change.org provided no information to readers that petitions could have been written with AI. However, if AI suspicion is driving petition outcomes, this problem might worsen as Change.org increases their advertising, and consequently readers’ suspicion, of the AI tool.
@@
Our third hypothesis for why petition outcomes were unchanged is not about how readers perceive others’ petitions, but how authors feel and act on their own petitions. Petition writers may feel diminished ownership over petitions they write with AI, and therefore be less likely to take actions that would attract signatures and comments, such as sharing the petition on social media. Previous literature on ownership in AI systems supports this explanation, showing that writers do not consider themselves to be the owners of text that is written with substantial AI support, especially when they use AI tools that require minimal user involvement [draxler2024ai, kadma2024inclusion, xu2024what, he2025which]. As perceptions of ownership contribute to how people construct their personal and collective identities [pierce2003state], and identity is central to sharing behaviors on social media [zhao2013faces], we would expect the diminished ownership to result in writers who are less likely to promote their petition in online communities. One way to address this concern could be through adding friction to the AI-assisted petition writing process, such as by requiring a lengthy user input or prompting the user to edit the draft produced by the AI tool. Research in platform design has demonstrated that friction can improve decision-making in collaborative human-AI systems by forcing users to have cognitive engagement [buccinca2021trust]. Such mechanisms could potentially improve authors’ perceptions of ownership and their likelihood of sharing petitions.
@@
It is especially important to consider the impact of AI integration on Change.org, where AI integration can influence which causes succeed and shift how users connect to their own communities. We found that AI has already caused a dramatic shift to petition lexical features on Change.org and driven platform-level homogenization. If AI continues to homogenize the language of protest, it has the capacity to undermine signals of petition authenticity and legitimacy [wojtowicz2025undermining]. Already, lexical features that were once highly predictive of petition outcomes have become weakened by platform-wide AI-driven shifts in lexical features. Regardless of lexical style, the introduction of in-platform AI can impact readers’ evaluations and writers’ engagement with their communities. Furthermore, with AI minimizing the cognitive load of petition-writing, petition writers might not only feel diminished ownership but also less engagement with the petition’s cause and their community at large. In sum, AI integration in petition-writing might significantly change the dynamics of civic engagement in ways that are difficult to predict and measure. Overall, our study serves as a cautionary tale for platforms that seek to integrate AI into their platforms: while user-generated content may change, AI tools may not yield dividends in participation or effectiveness, while potentially introducing additional harms.
@@
Change.org conducted a staggered roll-out of its AI writing tool by geographic region. The United States, Great Britain, and Canada were among the countries that participated in A/B testing during March-September 2023 and then received a full launch of the tool in October, 2023. However, there were several countries that had delayed access to the AI tool: Australia, New Zealand, and India. Other countries, like China, never received access. We used Australia as the control for our difference-in-differences analysis, as it showed adequately similar pre-AI trends in lexical features and outcomes in the United States, Great Britain, and Canada. While New Zealand, India, and China failed to demonstrate adequately similar pre-AI trends, we include petitions written in these countries in supplementary synthetic control analyses, included in the SI appendix.

### v3

@@ Abstract
references.bib] How does the integration of AI writing tools change content production, participation, and outcomes on online platforms? New AI writing tools and assistants based on Large Language Models (LLMs) are capable of generating high-quality text that is more articulate, persuasive, and creative than if the users had written the text themselves [Noy2023-sf, doshiHauser2024, matz2024potential, argyle2025testing, kumar2025human]. These tools and assistants are increasingly introduced into online platform user flows [experienceAmazonSellersCan2024, WriteAIFacebook]. However, the effect of in-platform AI writing tools on content and its outcomes on online platforms remains uncertain, with conflicting evidence regarding the contexts in which AI can augment human performance [vaccaroWhenCombinationsHumans2024]. While generally helpful, the effect of AI assistance on text performance could be task-specific; AI might not be as effective in tasks that are more nuanced or require personal context. Furthermore, AI writing tools might reduce the overall diversity of text produced across writers [doshiHauser2024, wu2024generativemonoculturelargelanguage, haoArtificialIntelligenceTools2026, agarwal2025ai, anderson2024homogenization], perhaps leading to less compelling, individualized text. To study the potential trade-offs of AI writing tools on a real-world online platform, we explore AI-assisted petition-writing on Change.org.
Change.org is a social advocacy platform where individuals around the world can write, read, and sign petitions aiming to effect change in their communities. Change.org aims to facilitate democratic deliberation and advocacy by enabling individuals to enact grassroots change and connect with their community around shared issues. By gathering community support in the form of signatures and comments, petitions are able to gain media attention and persuade decision-makers. In 2023, Change.org integrated an AI writing tool built on an OpenAI-provided LLM into the petition writing workflow. The site’s home page encouraged visitors to “use AI to create a compelling petition in minutes.” Petition writers can use this tool to generate petition drafts from only a short, user-entered description of a cause. Petition outcomes, and thereby the outcomes of the causes they champion, may be influenced by AI tools that impact the petitions’ content. Conceivably, integration of AI could bolster the effectiveness of the Change.org platform, as well as individual petitions, by lowering the barrier to participation and enabling users to produce high quality content with ease. As suggested by prior research, it is possible that this lower barrier for participation could increase petition production and completion rates on the platform, while improving the quality of content produced by low-skill writers [Noy2023-sf, brynjolfsson2025generativeAIatwork].
In this work, we leveraged a unique natural experiment to estimate the causal impact of access to the Change.org AI writing tool on published petitions in terms of lexical features and petition outcomes. Historically, the user experience on Change.org had been similar in the United States, Great Britain, Canada and Australia. However, this newly introduced AI tool was first rolled out in the United States, Great Britain, and Canada, with a delayed release in Australia. We conducted a difference-in-differences analysis that leverages this staggered roll-out in AI tool access to measure the causal impact. Unlike off-platform AI tools, the Change.org in-platform AI tool is uniquely tailored to the task and thereby channels AI usage through a single model, a single backend prompt, and standard interface. This in-platform AI tool and its staggered roll-out allowed us to avoid concerns that arise in many observational studies of generative AI use in the wild, where AI use can be difficult to detect and driven by idiosyncratic user flows.
Our research questions are twofold: (a) does access to the AI-generated draft tool result in changes to petition lexical features, and (b) does access to the tool result in changes to petition outcomes? To study these questions, we collected 1.5 million petitions written on Change.org between 2022 and 2024. We then estimated the impact of AI tool access using trends in petitions produced during the ten weeks of differential access between the full launch of the AI tool in the treated countries and introduction of the AI tool in Australia. We find that access to the AI-generated draft tool significantly changed petition lexical features. Specifically, the lexical features of petitions written with access to AI had greater resemblance to the those of pre-AI petitions that achieved strong outcomes. We also show that AI access resulted in increased homogeneity: petitions became more similar to each other. At the same time, access to the AI tool failed to improve petition outcome metrics such as hitting comment and signature thresholds. The shift in features and lack of improvement to petition outcomes is not driven by changes to platform participation; user participation remains constant after access to AI. In fact, we show that the observed trends of lexical features and outcomes are replicated when looking at repeat petition writers who wrote petitions with and without the AI tool access. Ultimately, our analyses show that integration of AI does not bolster the effectiveness of Change.org, despite significantly altering the characteristics of the petition text.
@@ Results
In our main analysis we use Australia as the control country to estimate causal effects. We confirm that our estimates are robust to multiple geographic regions with a synthetic control analysis that leverages additional control countries (New Zealand and India). The results of the synthetic control analysis are consistent with the difference-in-differences analysis, demonstrating that our causal estimates are consistent across regions and outside of Australia.
@@
We augmented our pre-registered analysis with a difference-in-differences study of the impact of AI access on platform homogeneity. The results show an increase in inter-petition homogeneity, as measured through the mean pairwise similarity between petitions [wu2024generativemonoculturelargelanguage, padmakumar2023does]. To measure homogeneity, we first created 3,072-dimension embeddings with the OpenAI text-embedding-3-large model [^3] https://platform.openai.com/docs/models/text-embedding-3-large for all petitions. We then calculated the cosine similarity between all pairs of petitions posted in a given week and country, and took the mean of these pairwise scores to obtain the mean pairwise similarity for that week and country. We used these weekly country-level time series of homogeneity to compare trends in treated and untreated countries through a dynamic difference-in-differences estimation.
@@
Petition Text Quality
We find that petitions written with access to the AI tool have, on average, higher writing quality and persuasiveness. One possible explanation for the lack of improvement to platform-level outcomes could be that the AI tool is simply not very effective: petitions produced with the tool could have lower quality text or be less persuasive than petitions written without the tool. To understand the quality of petitions written with and without AI access, we conducted a labeling task in which participants labeled petitions for writing quality and persuasiveness. Our descriptive results (full details in the SI Appendix) demonstrate that average petition writing quality is significantly higher among petitions written with in-platform AI access, relative to petitions written without AI access (t = 9.76, df = 412.46, p << 0.001, d = 0.96). Average petition persuasiveness scores are similarly significantly higher among petitions written with in-platform AI access (t = 8.75, df = 414.56, p << 0.001, d = 0.86). In the SI Appendix, we show that prior to AI access, writing quality and persuasiveness are highly associated with lexical features and predictive of petition outcomes. However, after AI access, writing quality and persuasiveness are weaker, or negative, predictors of petition outcomes. Overall, as writing quality and petition persuasiveness are significantly higher among petitions written with access to in-platform AI, it does not seem that changes to the quality of petition text explain the unimproved platform-level outcomes that we observe. We consider potential explanations and ramifications in the Discussion section. Next, we consider participation as another potential explanation for unimproved platform-level outcomes.
@@
To test the consistency of platform participation, we used pre-AI user participation data from treated and untreated countries to fit a Bayesian Structural Time Series (BSTS) model (see details in SI Appendix) [brodersen2015causalimpact]. This BSTS model uses pre-AI long-term trends and seasonal fluctuations to predict a synthetic control for counterfactual participation in treated countries had there been no AI tool launch. This forecast for platform participation closely reflected actual user participation, with actual participation falling within a 95% credible interval of forecast participation. Through comparing the synthetic counterfactual participation against actual platform participation, we do not find evidence of a significant effect of AI access on user participation, in either direction ($\Delta$=-0.43, 95% CI [-2.75, 1.95], p = 0.344).
@@
We further demonstrate that changes to platform composition were not responsible for changes to platform trends by showing that the platform-level trends—shifting petition features and unimproved outcomes—are consistent even when examining trends within repeat users. We focus our analysis here on one lexical feature (word count) and one outcome (comment engagement). The results for the other variables were similar, and are shown in the SI Appendix. We do not conduct a panel difference-in-differences analysis because of the rarity of repeat petition writing; only 7.6% of users in our dataset wrote multiple petitions, which amounted to too few untreated users for a reliable balanced panel. Instead, we model the predictive relationship between AI access and word count or comment engagement among repeat writers from treated countries (the United States, Great Britain, and Canada) over the full period for which we collected data.
@@ Discussion
Our work provides a causal analysis of the impact of AI integration into a real-world online platform. Our findings extend prior work on the impact of AI on online platforms at large: while online platforms might be influenced by both off-platform AI and in-platform AI tools, past studies have focused on general-use off-platform AI. As online platforms increasingly integrate AI, it is important to understand the potential impact on platform dynamics. In our case, the staggered roll-out of the Change.org in-platform AI tool allowed us to circumvent the challenges of noisy AI detection, and instead directly measure the impact of in-platform AI access through causal methods. We find that in-platform AI on Change.org shifts petition lexical features, likely improves quality, and increases site-wide homogeneity. At the same time, the in-platform AI tool did not improve platform outcomes, nor did it seem to affect platform participation.
Our results, and prior research, suggest that it may be the case that with AI, petition-writers could expend less effort to produce a petition of higher quality to a petition written without AI [Noy2023-sf]. Prior literature, as well as our results on petition writing quality and persuasiveness, suggests that the average quality of petition text and arguments would improve with AI assistance [HAGEN2016783, chen2019multi, schoenegger2025large, Noy2023-sf].
So, why doesn’t AI access improve platform-level outcomes on Change.org? We have three hypotheses for mechanisms that might drive the lack of change in outcomes: first, AI might not improve the substance of petitions; second, AI might increase readers’ AI suspicion; and third, AI-assisted writing might reduce petition-writer ownership and effort.
@@
First, even with higher text quality, petition outcomes might not improve if the petitions generated by the AI tool do not have improved substance. AI tools are not always guaranteed to augment human performance [vaccaroWhenCombinationsHumans2024]. While prior literature has explained human-AI collaboration through a “jagged technological frontier” in which AI either enhances human productivity and performance (inside the frontier) or performs worse than humans alone (outside the frontier) [dellacquaNavigatingJaggedTechnological2023], it may be the case that the technological capability of AI does not drive petition-writing success. In particular, the technological capabilities of AI to produce believable, high-quality text seem to be insufficient to explain the outcomes of content produced on this type of advocacy platform, where evaluations are determined by readers’ dynamic, subjective, and social perceptions. Specifically, it is possible that petitions, with their real-world, context-specific missions, fall outside of AI’s capabilities for augmentation because the task requires personal and community-specific knowledge and engagement [wojtowicz2025undermining]. For example, we show in the SI that petitions written with access to AI tended to use less specific language in the petition title. In sum, AI on Change.org might contribute to what others have called “semantic garbage” or “AI slop” [feuerriegel2023research, floridi2020gpt, moller2025impact, tullis2025sifting]: although AI shapes petitions to resemble previous articulate, well-researched petitions, the changes may be superficial.
@@
Our second proposed mechanism is AI suspicion: readers’ perceptions of petitions, and the Change.org platform, might deteriorate if they suspect the use of AI. Prior work has shown that AI suspicion has been attributed to social penalties since the release of ChatGPT in 2022 [jakesch2023human, reif2025socialpenalty]. While people use biased heuristics to detect AI [jakesch2023human], when people suspect AI use, they may perceive the content as less trustworthy and assign lower evaluations to the writers [kadoma2025generative, reif2025socialpenalty, jakesch2019ai, hohenstein2023artificial]. If the AI-written petitions induce higher reader suspicion, outcomes are indeed likely to be affected. This hypothesis is difficult to test with our data since human perceptions of AI assistance may have shifted considerably in the year since our study’s petitions were posted [duan2024understanding, matatov2024examining]. On the other hand, there are reasons to discount the “AI suspicion” hypothesis in this case. Petition readers were not primed to suspect AI use on Change.org; during the ten-week period for which we estimate the causal impact of AI, Change.org provided no information to readers that petitions could have been written with AI (we show the Change.org home page as of November 1, 2023 in SI Appendix). Despite the absence of platform-provided signals of AI use, readers may have heightened AI suspicion driven by the lexical style, paragraph structure, and vocabulary of petitions written with AI access. If AI suspicion is driving petition outcomes, this problem might worsen as Change.org increases their advertising, and consequently readers’ suspicion, of the AI tool.
@@
Our third hypothesis is author commitment: when writing with AI, authors may feel and act differently toward their own petitions. Petition writers may feel diminished ownership over petitions they write with AI, and therefore be less likely to take actions that would attract signatures and comments, such as sharing the petition on social media. Previous literature on ownership in AI systems supports this explanation, showing that writers do not consider themselves to be the owners of text that is written with substantial AI support, especially when they use AI tools that require minimal user involvement [draxler2024ai, kadma2024inclusion, xu2024what, he2025which]. As perceptions of ownership contribute to how people construct their personal and collective identities [pierce2003state], and identity is central to sharing behaviors on social media [zhao2013faces], we would expect the diminished ownership to result in writers who are less likely to promote their petition in online communities. As off-platform promotion is critical to petition success [prosser2011DoSocialNetworking, proskurnia2017PredictingtheSuccess, dumas2015ExaminingPoliticalMobilization, Carlson2019PleaseSignHere, Harrison2022ExploringEpetitioning], with less promotion petitions would likely get fewer comments and signatures. One way to address this concern could be through adding friction to the AI-assisted petition writing process, such as by requiring a lengthy user input or prompting the user to edit the draft produced by the AI tool. Research in platform design has demonstrated that friction can improve decision-making in collaborative human-AI systems by forcing users to have cognitive engagement [buccinca2021trust]. Such mechanisms could potentially improve authors’ perceptions of ownership and their likelihood of sharing petitions.
@@
It is especially important to consider the impact of AI integration on Change.org, where AI integration can influence which causes succeed and shift how users connect to their own communities. We found that AI has already caused a dramatic shift to petition lexical features on Change.org and driven platform-level homogenization. If AI continues to homogenize the language of protest, it has the capacity to undermine signals of petition authenticity and legitimacy [wojtowicz2025undermining]. As we show in the SI Appendix, the predictive strength of features of petition text (i.e., lexical features, writing quality, persuasiveness) has weakened, or reversed, with platform-wide shifts that correspond with in-platform AI access. Regardless of lexical style, the introduction of in-platform AI can impact readers’ evaluations and writers’ engagement with their communities. Furthermore, with AI minimizing the cognitive load of petition-writing, petition writers might not only feel diminished ownership but also less engagement with the petition’s cause and their community at large. In sum, AI integration in petition-writing might significantly change the dynamics of civic engagement in ways that are difficult to predict and measure. Despite potential harms, Change.org has deepened the integration of the AI tool into the petition writing flow. As of early 2026, some users visiting the Change.org home page are no longer offered the choice to start a petition without AI support; instead, they are met with a text-box that automatically launches the AI tool (see example in SI Appendix). Overall, our study serves as a cautionary tale for platforms that seek to integrate AI into their platforms: while user-generated content may change, AI tools may not yield dividends in participation or effectiveness, while potentially introducing additional harms.
@@
Prior research into online petitions has demonstrated that e-petitions are dependent on off-platform promotion through social networks, news outlets, or direct shares to gain signatures [prosser2011DoSocialNetworking, proskurnia2017PredictingtheSuccess, dumas2015ExaminingPoliticalMobilization, Carlson2019PleaseSignHere, Harrison2022ExploringEpetitioning]. Crowdfunding platforms, which operate similarly to Change.org, have been found to display “friendfunding”, by which campaigns tend to be funded by friends and family, unless shared broadly on external social media platforms [lee2022NewDigitalSafetyNet, borst2018FromFriendfunding]. Based on this extensive prior work, we hypothesize that few readers browse Change.org in pursuit of petitions to support, but instead find petitions off-platform through social networks, news coverage, or direct shares from friends.
@@
Change.org conducted a staggered roll-out of its AI writing tool by geographic region. The United States, Great Britain, and Canada were among the countries that participated in A/B testing during March-September 2023 and then received a full launch of the tool in October, 2023. However, there were several countries that had delayed access to the AI tool: Australia, New Zealand, and India. Other countries, like China, never received access. We used Australia as the control for our difference-in-differences analysis, as it showed adequately similar pre-AI trends in lexical features and outcomes in the United States, Great Britain, and Canada. While New Zealand and India failed to demonstrate adequately similar pre-AI trends, we include petitions written in these countries in supplementary synthetic control analyses, included in the SI appendix.